OpenStudy (anonymous):

simplifying this radical √12x3y8

zepdrix (zepdrix):

\[\huge \sqrt{\color{#3366CF}{12}\color{#3399AA}{x^3}\color{#662FFF}{y^8}}=\sqrt{\color{#3366CF}{12}}\cdot \sqrt{\color{#3399AA}{x^3}}\cdot \sqrt{\color{#662FFF}{y^8}}\]Hmm so if we break it up, we can deal with each piece separately. Let's look at the 12 first.

OpenStudy (anonymous):

6X2= 12

zepdrix (zepdrix):

We want a FACTOR of 12 that is a perfect square. Perfects squares are as such, 4, 9, 16, 25, 36, ...\[\large \sqrt4=2\]See how we can take the square root and get a nice even number? Hmm Those are factors of 12 yes, but neither 6 nor 2 is a perfect square.

zepdrix (zepdrix):

Can you think of some other factors of 12? c:

OpenStudy (anonymous):

2X5 / 4X3

zepdrix (zepdrix):

4x3, yah let's try that c: \[\huge \sqrt{\color{#3366CF}{12}}\qquad =\qquad \sqrt{\color{#3366CF}{4}}\cdot \sqrt{\color{#3366CF}{3}}\qquad =\qquad \color{#3366CF}{2}\sqrt{\color{#3366CF}{3}}\]

zepdrix (zepdrix):

Understand what we did there? :o

OpenStudy (anonymous):

mhmm 2X2=4 and nothing goes into 3 ....so yeah i undertstand

zepdrix (zepdrix):

That simplifies the 12, that's as far as we can break it down. So we're 1/3 done c:

zepdrix (zepdrix):

For the X term, you'll want to remember how to rewrite a radical using FRACTIONS in the exponent. Here is how our X term would change,\[\huge \sqrt{\color{#3399AA}{x^3}} \qquad = \qquad (\color{#3399AA}{x^3})^{1/2}\]

OpenStudy (anonymous):

my teacher was talking about that but i didnt get it

zepdrix (zepdrix):

\[\huge \sqrt{\color{#3399AA}{x^3}} \qquad = \qquad (\color{#3399AA}{x^3})^{1/2}\qquad = \qquad \color{#3399AA}{x^{3/2}}\]

zepdrix (zepdrix):

Hmm ok lemme see if I can think of a simple example :O

zepdrix (zepdrix):